# Orchestrated Robust Controller for the Precision Control of Heavy-duty Hydraulic Manipulators

### v2

@@ VIII Conclusion
$\begin{split}\nu_{a}&=\left(f_{ppr}-f_{pp}\right)^{2}/(2\,\beta\,k_{xp})+\sum_{j=1,2}\left(f_{pjr}-f_{pj}\right)^{2}/(2\,\beta\,k_{xj})\\ &+\sum_{i=1,2,3}\left(f_{pwir}-f_{pwi}\right)^{2}/(2\,\beta\,k_{xwi})+\frac{1}{2\gamma_{dp}}\left(\tilde{\theta}_{dp}^{T}\,\tilde{\theta}_{dp}\right)\\ &+\sum_{j=1,2}\frac{1}{2\gamma_{dj}}\left(\tilde{\theta}_{dj}^{T}\,\tilde{\theta}_{dj}\right)+\sum_{i=1,2,3}\frac{1}{2\gamma_{dwi}}\left(\tilde{\theta}_{dwi}^{T}\,\tilde{\theta}_{dwi}\right)\\ &+\frac{1}{2\gamma_{vp}}\left(\tilde{\theta}_{vp}^{T}\,\tilde{\theta}_{vp}\right)+\sum_{j=1,2}\frac{1}{2\gamma_{vj}}\left(\tilde{\theta}_{vj}^{T}\,\tilde{\theta}_{vj}\right)\\ &+\sum_{i=1,2,3}\frac{1}{2\gamma_{vwi}}\left(\tilde{\theta}_{vwi}^{T}\,\tilde{\theta}_{vwi}\right)+\sum_{j=1,2}\frac{1}{2\gamma_{fj}}\left(\tilde{\theta}_{fj}^{T}\,\tilde{\theta}_{fj}\right)\\ &+\frac{1}{2\gamma_{fp}}\left(\tilde{\theta}_{fp}^{T}\,\tilde{\theta}_{fp}\right)+\sum_{i=1,2,3}\frac{1}{2\gamma_{fwi}}\left(\tilde{\theta}_{fwi}^{T}\,\tilde{\theta}_{fwi}\right)\\ &+\frac{1}{2\delta_{p}}\left(\tilde{W}_{ap}^{T}\,\tilde{W}_{ap}\right)+\sum_{j=1,2}\frac{1}{2\delta_{j}}\left(\tilde{W}_{aj}^{T}\,\tilde{W}_{aj}\right)\\ &+\sum_{i=1,2,3}\frac{1}{2\delta_{wi}}\left(\tilde{W}_{awi}^{T}\,\tilde{W}_{awi}\right)+\sum_{i=1,2,3}\frac{1}{2\bar{\delta}_{wi}}\left(\tilde{\varepsilon}_{awi}\right)^{2}\\ &+\frac{1}{2\bar{\delta}_{p}}\left(\tilde{\varepsilon}_{ap}\right)^{2}+\sum_{j=1,2}\frac{1}{2\bar{\delta}_{j}}\left(\tilde{\varepsilon}_{aj}\right)^{2}+\frac{1}{2\delta_{p\theta}}\left(\tilde{\theta}_{p}^{T}\tilde{\theta}_{p}\right)\\ &+\sum_{j=1,2}\frac{1}{2\delta}_{j\theta}\left(\tilde{\theta}_{j}^{T}\,\tilde{\theta}_{j}\right)+\sum_{i=1,2,3}\frac{1}{2\delta}_{wi\theta}\left(\tilde{\theta}_{awi}^{T}\,\tilde{\theta}_{awi}\right)\end{split}$ (130)
@@
$\begin{split}\dot{\nu}_{a}&\leq-\frac{k_{fp}}{k_{xp}}\,(f_{ppr}-f_{pp})^{2}-\sum_{j=1,2}\frac{k_{fj}}{k_{xj}}(f_{pjr}-f_{pj})^{2}\\ &-\sum_{i=1,2,3}\frac{k_{fwi}}{k_{xwi}}(f_{pwir}-f_{pwi})^{2}+\,VPFs\\ &-0.5\gamma_{dp0}\tilde{\theta}_{dp}^{T}\,\tilde{\theta}_{dp}-0.5\sum_{j=1,2}\gamma_{dj0}\,\tilde{\theta}_{dj}^{T}\tilde{\theta}_{dj}\\ &-0.5\sum_{i=1,2,3}\gamma_{dwi0}\,\tilde{\theta}_{dwi}^{T}\tilde{\theta}_{dwi}-0.5\gamma_{vp0}\tilde{\theta}_{vp}^{T}\tilde{\theta}_{vp}\\ &-0.5\sum_{j=1,2}\gamma_{vj0}\tilde{\theta}_{vj}^{T}\tilde{\theta}_{vj}-0.5\sum_{i=1,2,3}\gamma_{vwi0}\tilde{\theta}_{vwi}^{T}\tilde{\theta}_{vwi}\\ &-0.5{\gamma_{fp0}}\tilde{\theta}_{fp}^{T}\tilde{\theta}_{fp}-0.5\sum_{i=1,2,3}{\gamma_{fwi0}}\tilde{\theta}_{fwi}^{T}\tilde{\theta}_{fwi}\\ &-0.5\sum_{j=1,2}{\gamma_{fj0}}\tilde{\theta}_{fj}^{T}\tilde{\theta}_{fj}-0.5\sum_{i=1,2,3}{\delta_{wi0}}\tilde{W}_{awi}^{T}\tilde{W}_{awi}\\ &-0.5{\delta_{p0}}\tilde{W}_{ap}^{T}\tilde{W}_{ap}-0.5\sum_{j=1,2}{\delta_{j0}}\tilde{W}_{aj}^{T}\tilde{W}_{aj}\\ &-0.5{\bar{\delta}_{p0}}\tilde{\varepsilon}_{ap}\tilde{\varepsilon}_{ap}-0.5\sum_{j=1,2}{\bar{\delta}_{j0}}\tilde{\varepsilon}_{aj}\tilde{\varepsilon}_{aj}\\ &-0.5\sum_{i=1,2,3}{\bar{\delta}_{wi0}}\tilde{\varepsilon}_{awi}\tilde{\varepsilon}_{awi}-0.5{\delta_{p\theta 0}}\tilde{\theta}_{p}^{T}\tilde{\theta}_{p}\\ &-0.5\sum_{j=1,2}{\delta_{j\theta 0}}\tilde{\theta}_{j}^{T}\tilde{\theta}_{j}-0.5\sum_{i=1,2,3}{\delta_{wi\theta 0}}\tilde{\theta}_{wi}^{T}\tilde{\theta}_{wi}\\ &+0.5\gamma_{dp0}\theta_{dp}^{T}\,\theta_{dp}+0.5\sum_{j=1,2}\gamma_{dj0}\,\theta_{dj}^{T}\theta_{dj}\\ &+0.5\sum_{i=1,2,3}\gamma_{dwi0}\,\theta_{dwi}^{T}\theta_{dwi}+0.5\gamma_{vp0}\theta_{vp}^{T}\theta_{vp}\\ &+0.5\sum_{j=1,2}\gamma_{vj0}\theta_{vj}^{T}\theta_{vj}+0.5\sum_{j=1,2}{\bar{\delta}_{j0}}{\varepsilon}_{aj}{\varepsilon}_{aj}\\ &+0.5\sum_{i=1,2,3}\gamma_{vwi0}\theta_{vwi}^{T}\theta_{vwi}+0.5{\gamma_{fp0}}\theta_{fp}^{T}\theta_{fp}\\ &+0.5\sum_{j=1,2}{\gamma_{fj0}}\theta_{fj}^{T}\theta_{fj}+0.5\sum_{i=1,2,3}{\gamma_{fwi0}}\theta_{fwi}^{T}\theta_{fwi}\\ &+0.5{\delta_{p0}}{W}_{ap}^{T}{W}_{ap}+0.5\sum_{j=1,2}{\delta_{j0}}{W}_{aj}^{T}{W}_{aj}\\ &+0.5\sum_{i=1,2,3}{\delta_{wi0}}{W}_{awi}^{T}{W}_{awi}+0.5{\bar{\delta}_{p0}}{\varepsilon}_{ap}{\varepsilon}_{ap}\\ &+0.5\sum_{i=1,2,3}{\bar{\delta}_{wi0}}{\varepsilon}_{awi}{\varepsilon}_{awi}+0.5{\delta_{p\theta 0}}\theta_{p}^{T}\theta_{p}\\ &+0.5\sum_{j=1,2}{\delta_{j\theta 0}}\tilde{\theta}_{j}^{T}\theta_{j}+0.5\sum_{i=1,2,3}{\delta_{wi\theta 0}}\theta_{wi}^{T}\theta_{wi}\\ &\leq-\mu_{a}\nu_{a2}+\mu_{a0}+VPFs.\end{split}$ (134)

### v4

@@ IX Conclusion
$\begin{split}\nu_{a}&=\left(f_{ppr}-f_{pp}\right)^{2}/(2\,\beta\,k_{xp})+\sum_{j=1,2}\left(f_{pjr}-f_{pj}\right)^{2}/(2\,\beta\,k_{xj})\\ &+\sum_{i=1,2,3}\left(f_{pwir}-f_{pwi}\right)^{2}/(2\,\beta\,k_{xwi})+\frac{1}{2\gamma_{dp}}\left(\tilde{\theta}_{dp}^{T}\,\tilde{\theta}_{dp}\right)\\ &+\sum_{j=1,2}\frac{1}{2\gamma_{dj}}\left(\tilde{\theta}_{dj}^{T}\,\tilde{\theta}_{dj}\right)+\sum_{i=1,2,3}\frac{1}{2\gamma_{dwi}}\left(\tilde{\theta}_{dwi}^{T}\,\tilde{\theta}_{dwi}\right)\\ &+\frac{1}{2\gamma_{vp}}\left(\tilde{\theta}_{vp}^{T}\,\tilde{\theta}_{vp}\right)+\sum_{j=1,2}\frac{1}{2\gamma_{vj}}\left(\tilde{\theta}_{vj}^{T}\,\tilde{\theta}_{vj}\right)\\ &+\sum_{i=1,2,3}\frac{1}{2\gamma_{vwi}}\left(\tilde{\theta}_{vwi}^{T}\,\tilde{\theta}_{vwi}\right)+\sum_{j=1,2}\frac{1}{2\gamma_{fj}}\left(\tilde{\theta}_{fj}^{T}\,\tilde{\theta}_{fj}\right)\\ &+\frac{1}{2\gamma_{fp}}\left(\tilde{\theta}_{fp}^{T}\,\tilde{\theta}_{fp}\right)+\sum_{i=1,2,3}\frac{1}{2\gamma_{fwi}}\left(\tilde{\theta}_{fwi}^{T}\,\tilde{\theta}_{fwi}\right)\\ \end{split}$
$\begin{split}&+\frac{1}{2\delta_{p}}\left(\tilde{W}_{ap}^{T}\,\tilde{W}_{ap}\right)+\sum_{j=1,2}\frac{1}{2\delta_{j}}\left(\tilde{W}_{aj}^{T}\,\tilde{W}_{aj}\right)\\ &+\sum_{i=1,2,3}\frac{1}{2\delta_{wi}}\left(\tilde{W}_{awi}^{T}\,\tilde{W}_{awi}\right)+\sum_{i=1,2,3}\frac{1}{2\bar{\delta}_{wi}}\left(\tilde{\varepsilon}_{awi}\right)^{2}\\ &+\frac{1}{2\bar{\delta}_{p}}\left(\tilde{\varepsilon}_{ap}\right)^{2}+\sum_{j=1,2}\frac{1}{2\bar{\delta}_{j}}\left(\tilde{\varepsilon}_{aj}\right)^{2}+\frac{1}{2\delta_{p\theta}}\left(\tilde{\theta}_{p}^{T}\tilde{\theta}_{p}\right)\\ &+\sum_{j=1,2}\frac{1}{2\delta}_{j\theta}\left(\tilde{\theta}_{j}^{T}\,\tilde{\theta}_{j}\right)+\sum_{i=1,2,3}\frac{1}{2\delta}_{wi\theta}\left(\tilde{\theta}_{awi}^{T}\,\tilde{\theta}_{awi}\right)\end{split}$ (130)
@@
$\begin{split}\dot{\nu}_{a}&\leq-\frac{k_{fp}}{k_{xp}}\,(f_{ppr}-f_{pp})^{2}-\sum_{j=1,2}\frac{k_{fj}}{k_{xj}}(f_{pjr}-f_{pj})^{2}\\ &-\sum_{i=1,2,3}\frac{k_{fwi}}{k_{xwi}}(f_{pwir}-f_{pwi})^{2}+\,VPFs\\ &-0.5\gamma_{dp0}\tilde{\theta}_{dp}^{T}\,\tilde{\theta}_{dp}-0.5\sum_{j=1,2}\gamma_{dj0}\,\tilde{\theta}_{dj}^{T}\tilde{\theta}_{dj}\\ &-0.5\sum_{i=1,2,3}\gamma_{dwi0}\,\tilde{\theta}_{dwi}^{T}\tilde{\theta}_{dwi}-0.5\gamma_{vp0}\tilde{\theta}_{vp}^{T}\tilde{\theta}_{vp}\\ &-0.5\sum_{j=1,2}\gamma_{vj0}\tilde{\theta}_{vj}^{T}\tilde{\theta}_{vj}-0.5\sum_{i=1,2,3}\gamma_{vwi0}\tilde{\theta}_{vwi}^{T}\tilde{\theta}_{vwi}\\ &-0.5{\gamma_{fp0}}\tilde{\theta}_{fp}^{T}\tilde{\theta}_{fp}-0.5\sum_{i=1,2,3}{\gamma_{fwi0}}\tilde{\theta}_{fwi}^{T}\tilde{\theta}_{fwi}\\ &-0.5\sum_{j=1,2}{\gamma_{fj0}}\tilde{\theta}_{fj}^{T}\tilde{\theta}_{fj}-0.5\sum_{i=1,2,3}{\delta_{wi0}}\tilde{W}_{awi}^{T}\tilde{W}_{awi}\\ &-0.5{\delta_{p0}}\tilde{W}_{ap}^{T}\tilde{W}_{ap}-0.5\sum_{j=1,2}{\delta_{j0}}\tilde{W}_{aj}^{T}\tilde{W}_{aj}\\ &-0.5{\bar{\delta}_{p0}}\tilde{\varepsilon}_{ap}\tilde{\varepsilon}_{ap}-0.5\sum_{j=1,2}{\bar{\delta}_{j0}}\tilde{\varepsilon}_{aj}\tilde{\varepsilon}_{aj}\\ &-0.5\sum_{i=1,2,3}{\bar{\delta}_{wi0}}\tilde{\varepsilon}_{awi}\tilde{\varepsilon}_{awi}-0.5{\delta_{p\theta 0}}\tilde{\theta}_{p}^{T}\tilde{\theta}_{p}\\ &-0.5\sum_{j=1,2}{\delta_{j\theta 0}}\tilde{\theta}_{j}^{T}\tilde{\theta}_{j}-0.5\sum_{i=1,2,3}{\delta_{wi\theta 0}}\tilde{\theta}_{wi}^{T}\tilde{\theta}_{wi}\\ &+0.5\gamma_{dp0}\theta_{dp}^{T}\,\theta_{dp}+0.5\sum_{j=1,2}\gamma_{dj0}\,\theta_{dj}^{T}\theta_{dj}\\ &+0.5\sum_{i=1,2,3}\gamma_{dwi0}\,\theta_{dwi}^{T}\theta_{dwi}+0.5\gamma_{vp0}\theta_{vp}^{T}\theta_{vp}\\ &+0.5\sum_{j=1,2}\gamma_{vj0}\theta_{vj}^{T}\theta_{vj}+0.5\sum_{j=1,2}{\bar{\delta}_{j0}}{\varepsilon}_{aj}{\varepsilon}_{aj}\\ &+0.5\sum_{i=1,2,3}\gamma_{vwi0}\theta_{vwi}^{T}\theta_{vwi}+0.5{\gamma_{fp0}}\theta_{fp}^{T}\theta_{fp}\\ &+0.5\sum_{j=1,2}{\gamma_{fj0}}\theta_{fj}^{T}\theta_{fj}+0.5\sum_{i=1,2,3}{\gamma_{fwi0}}\theta_{fwi}^{T}\theta_{fwi}\\ &+0.5{\delta_{p0}}{W}_{ap}^{T}{W}_{ap}+0.5\sum_{j=1,2}{\delta_{j0}}{W}_{aj}^{T}{W}_{aj}\\ &+0.5\sum_{i=1,2,3}{\delta_{wi0}}{W}_{awi}^{T}{W}_{awi}+0.5{\bar{\delta}_{p0}}{\varepsilon}_{ap}{\varepsilon}_{ap}\\ &+0.5\sum_{i=1,2,3}{\bar{\delta}_{wi0}}{\varepsilon}_{awi}{\varepsilon}_{awi}+0.5{\delta_{p\theta 0}}\theta_{p}^{T}\theta_{p}\\ \end{split}$
$\begin{split}&+0.5\sum_{j=1,2}{\delta_{j\theta 0}}\tilde{\theta}_{j}^{T}\theta_{j}+0.5\sum_{i=1,2,3}{\delta_{wi\theta 0}}\theta_{wi}^{T}\theta_{wi}\\ &\leq-\mu_{a}\nu_{a2}+\mu_{a0}+VPFs.\end{split}$ (134)
@@
-D Unique Matrix Definition
The unique symmetric and positive-definite matrix in (105) can be defined as,
${}^{A}\mathcal{I}(\mathbf{s})=\begin{bmatrix}\mathbf{s}_{6}+\mathbf{s}_{7}&-0.5\mathbf{s}_{8}&-0.5\mathbf{s}_{10}&0.5\mathbf{s}_{2}\\ -0.5\mathbf{s}_{8}&\mathbf{s}_{5}+\mathbf{s}_{7}&-0.5\mathbf{s}_{9}&0.5\mathbf{s}_{3}\\ -0.5\mathbf{s}_{10}&-0.5\mathbf{s}_{9}&\mathbf{s}_{5}+\mathbf{s}_{6}&0.5\mathbf{s}_{4}\\ 0.5\mathbf{s}_{2}&0.5\mathbf{s}_{3}&0.5\mathbf{s}_{4}&\mathbf{s}_{1}\end{bmatrix}.$